OpenStudy (anonymous):

Fifty-Five perfect of the 4000 customers placed orders, using the company's online service. How many customers did not order online?

OpenStudy (anonymous):

help a brother out

OpenStudy (magbak):

I got this

OpenStudy (magbak):

i posted the ratio box if you need it

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

i just need the proportion for this one

OpenStudy (magbak):

It is 2200

OpenStudy (magbak):

Oh sorry the proportion is (400 x .55) there is no proportion

OpenStudy (magbak):

oh actually there is

OpenStudy (magbak):

it is 22--/55=x/45

OpenStudy (magbak):

i ment 2200/55= X/45

OpenStudy (magbak):

You get it

OpenStudy (anonymous):

yup thx!